OpenStudy (anonymous):

You are standing next to a well, and you have two jugs. One jug has a volume of 3 litres and the other one has a volume of 5 litres. Immersing either jug in the well will completely fill it. You want exactly 4 litres. How do you get that?

OpenStudy (anonymous):

There are two solutions that I know of by the way.

OpenStudy (anonymous):

You want to calculate what are called the Bezout Coefficients for 3 and 5. Then multiply them by 4.

OpenStudy (anonymous):

okay got it first fill the 5 then from 5 to 3 then you know there is 2 in 5 then from 5 to 3 and then 2 is in 3 then fill 5 again and full 3 ......4 in 5

OpenStudy (chaise):

Fill the 5 litre jug and then pour the water carefully into the 3 litre jug till it is full. You now have 2 litres left in the 5 litre jug. Now dump out the 3 litre jug and pour what's left in the 5 litre jug into the 3 litre jug. Now refill the 5 litre jug and fill up the 3 litre jug - should only take 1 litre to do this. What's left in the 5 litre jug will be 4 litres - just what you were asked for.

OpenStudy (anonymous):

you copy paste chaise

OpenStudy (chaise):

Oh, I wasn't allowed to copy and paste?

OpenStudy (anonymous):

Using Bezout's Identity, the gcd of 5 and 3 is 1, so there exist some integers x and y such that: 3x + 5y = 1 Those integers are x = 2, y = -1 multiplying by 4 gives: 3(4*2)+ 5*( 4*-1) = 4 so 3(8)+5(-4) = 4 you would fill up 8 times with the 3 liter jug, then take away 4 of the 5 liter jug.

OpenStudy (anonymous):

Surely that only works if you have a third jug big enough to accommodate at least 7 litres? Either way, that's a nice answer, so points!

OpenStudy (anonymous):

oh i didnt realize we were trying to accomplish this with only the two jugs at hand, my bad.

OpenStudy (anonymous):

ive seen questions before where they ask something like, "is it possible to measure 1 liter of water if you have a container that holds x liters, and one that holds y liters?" i guess i automatically assumed it was the same type.